OpenStudy (anonymous):

(a) find the slope of the tangent to the parabola y=3x^2+2x at the point whose x-coordinate is a. (b) at what point on the parabola is the tangent line parallel to the line y=10x-2? thanks!

OpenStudy (diamondboy):

can u use derivatives well?

OpenStudy (anonymous):

i don't know derivatives yet. i just start doing calculus on limits

OpenStudy (anonymous):

i got part a), that answer is 6a+2, but I'm not sure how to do part b)

OpenStudy (freckles):

so you found y'=6x+2 at y' at x=a you have 6a+2 so the slope of the tangent line at x=a is 6a+2 - So question b tells you to find at what point on the parabola the tangent line is paralel to y=10x-2 We know parallel lines have the same slope (and different y-intercept) so what is the slope of y=10x-2?

OpenStudy (freckles):

And you will set 6a+2=slope of(10x-2) and solve for a

OpenStudy (anonymous):

a=5/3x-2/3?

OpenStudy (freckles):

the slope of 10x-2 is?

OpenStudy (freckles):

\[6a+2=\text{ slope of }(10x+2) \\ 6a+2=10\]

OpenStudy (freckles):

10x-2 whatever

OpenStudy (freckles):

but still slope of 10x-2 is 10

OpenStudy (freckles):

@esam2 Do you understand that the slope of the tangent line has to be the same as the other line if we want them to be parallel?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a=4/3

OpenStudy (freckles):

slope of tangent line at x=a is 6a+2 slope of the other line is 10 we want to find when those slopes are the same 6a+2 is the same as 10 when a=? 6a=10-2 6a=8 a=8/6=4/3 yep now it says to find the point you found the x value such that the slopes are the same now find the y on the parabola for that given x

OpenStudy (freckles):

f(x)=3x^2+2x so f(4/3)=?

OpenStudy (anonymous):

is it 8? so it will be (4/3,8)?

OpenStudy (freckles):

\[f(\frac{4}{3})=3(\frac{4}{3})^2+2(\frac{4}{3}) \\ 3 \cdot \frac{4}{3} \cdot \frac{4}{3}+\frac{8}{3} \\ \frac{16}{3}+\frac{8}{3}=\frac{24}{3}=8\] yeah

OpenStudy (anonymous):

thanks!

OpenStudy (freckles):

np